Shift report 20000715 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Jul 15 2000 - 00:01:41 EDT

  • Next message: http account: "Shift report 20000715 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Jul 15 00:01:40 2000, Shift: 16:00-24:00
    Supervisor: sjs, with: sjs,am
    Log book#3 pages 7 - 7
    
    Shift Summary:
    ------------------------------------------------------------
    No beam:  "RHIC injection and ramping continues"
    
    Fixed Si pulser to allow for simultaneous 
    pulser and pedistal
    measurements.  Updated description of Si strip 
    detector arrangment and operation 
    in "Experiment Information" folder. The
    SuperMonitor program has been updated for the
    current arrangment, but the changes need to be committed
    to the CVS (Hiro will do this).  It should be noted that
    for each Si detector we are reading out 7 energy channels
    (for the seven strips) AND one temperature channel.  The
    temperature channel is a DC level and will not show a 
    pedistal.  Also, we are missing some connectors that are 
    necessary to add  pulser signal to all of the Si 
    channels.  At the moment, only 10-11 detectors can have a
    pulser signals applied at a given time.  We are also short
    one FERA for the Si detector readout.  The connector and 
    FERA shortages should be fixed next week once Ken gets back.   
    
    Replaced 10 bit ADC for the ZDC with an LCR2249W module 
    reclaimed from Kansas.   
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Jul 15 2000 - 00:05:06 EDT